OpenStudy (anonymous):

In a 30°- 60°- 90° right triangle, the longer leg is radical 3 cm. How long are the other two sides of the triangle? In a 30°- 60°- 90° right triangle, the longer leg is radical 3 cm. How long are the other two sides of the triangle? @Mathematics

OpenStudy (anonymous):

radical 1 and radical 2

OpenStudy (anonymous):

i think, let me check

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

i can give u the choices i have if that wud help?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

shorter leg 6 hypotenuse shorter leg 3 hypotenuse 6 shorter leg 3 hypotenuse shorter leg 6 hypotenuse 3

OpenStudy (anonymous):

whops read the question wrong

OpenStudy (anonymous):

lozl its ok :D

OpenStudy (anonymous):

y is their no number next to the hypothenus

OpenStudy (anonymous):

oh sorry its 6 radical 3

OpenStudy (anonymous):

the 2 of them that are blank are 6radical3

OpenStudy (anonymous):

i really have no idea what im doing sorry

OpenStudy (anonymous):

oh its fine :) i dont either

OpenStudy (anonymous):

i hate radicals,

OpenStudy (anonymous):

i feel ya

OpenStudy (anonymous):

or anything that looks like it log, squareroots that stuff makes feel feel sick

OpenStudy (anonymous):

I solved it using radical 3 as the longer leg of the triangle. Try solving it again with the right measurement. If you draw the triangle you'll be able to notice that the longer leg of it is determined by the 60º angle. Using the sin rule a/sinA = b/sinB = c/sinC 3^(1/2)/sin(60) -> 3^(1/2)/3^(1/2)/2 -> 1/2 so hip/sin(90) = 1/2 and smaller leg/sin(30) = 1/2 hip=1/2(sin(90)) smallerleg = 1/2(sin(30)) hip = 1/2 smaller leg = 1/4